Write K G equation
1
Expert's answer
2020-10-15T03:07:53-0400

Solution

The Klein–Gordon equation (K G equation) is related to the Schrödinger equation and also a relativistic wave equation.

for "\hbar=c=1" K G equation is given by

"\nabla^2\psi-\frac{ d^2\psi}{dt^2}=m^2\psi"